OpenStudy (anonymous):

a rectangular garden measures 5m by 7m. both dimensions are to be extended by the same amount so that the area of the garden is doubled . by how much should the dimensions increase, to the nearest tenth of a metre?

OpenStudy (anonymous):

You have 5 * 7= 35 and you want to multiply each side by the same number in order to get double area (70) so (5*a) * (7*a) = 70 where a is the number you want to multiply \[35a^2=70\] \[a^2=2\] \[a=\sqrt{2}\]